OpenStudy (anonymous):

Factor the polynomial. 42w10 + 24w6 Help! =) Please. . .

OpenStudy (anonymous):

We were not done yet.

OpenStudy (anonymous):

Once you find the GCF which is 6w^6 you divide that out of both terms

OpenStudy (anonymous):

6w^6(7w^4 + 4)

OpenStudy (anonymous):

You divide numerical numbers i.e. 42/6 = 7 and 24/6 = 4 and you subtract exponents 10 - 6 = w^4 and 6 - 6 = 0 so no w's

OpenStudy (anonymous):

Ohh, Okay Got It thanks!